# Infinity Synthetic Environments for Agentic Reinforcement Learning

### v2

@@ 6.2 Analysis on Verification Design
Table 6 compares three verification strategies. LLM-only verification relies on trajectory without being grounded in database state changes, which makes the reward signal unreliable. Consequently, it yields the weakest performance. Replacing the judge with code-only verification improves over LLM-only, but it can be brittle: when environment imperfections occur, rigid checks may incorrectly assign false negatives, as discussed in Sec. 3.3.1. In contrast, our code-augmented strategy combines structured verification signals (from state diffs and rule-based checks) with an advanced reasoning LLM, allowing the judge to use grounded evidence while remaining tolerant to imperfect environment signals. This hybrid design consistently achieves the best results across model scales and benchmarks, indicating that it provides a more robust reward for RL training in imperfect synthetic environments. Finally, the extra cost of invoking GPT-5 as the judge is about \$1.80 on average per training step (at most 1,024 samples), and the asynchronous setting makes augmented judge introduce negligible latency. We also study the format correctness reward in Appendix B.1.
@@ 6.3 Analysis on History-Aware Training
Table 7 compares our history-aware training against a variant using full history (w/o HL), under aligned and misaligned inference settings. When aligned, AWM w/ HL achieves the best results among all settings, indicating benefits from optimizing directly under truncated histories. In contrast, the full-history variant is relatively insensitive to misalignment. When truncated at inference, it even shows a slight improvement in $\tau^{2}$, consistent with truncation suppressing interference from earlier irrelevant turns. We believe history management should be treated as part of policy optimization rather than a purely inference-time heuristic.
@@ 6.4 Environment Scaling Curve
Figure 4 shows how agent scales with the number of training environments. 10 environments lead to severe performance degradation across all benchmarks, likely due to overfitting to the limited environment distribution. Scaling to 100 environments yields substantial gains, and further scaling to 526 environments continues to improve performance. This monotonic improvement highlights the importance of environment diversity for agentic RL. Due to compute constraints, we train on 526 environments rather than the full synthesized set. However, the aforementioned analysis confirms that diversity remains stable as the environment pool expands, suggesting that AWM can support scaling well beyond 1,000 environments with sustained benefits.
Figure 4: Scaling of AWM over environment sizes with 4B model.
@@ 7 Conclusion
In this paper, we propose Agent World Model, a scalable pipeline that synthesizes executable agentic environments by mirroring practical software development. Using AWM, we successfully scale to 1,000 environments with 35,062 tools and 10,000 tasks. These environments are code-driven and backed by SQL databases exposed through a unified MCP interface, supporting parallel isolated instances for large-scale agentic RL. Experiments on three benchmarks show that agents trained on our synthetic environments generalize well to out-of-distribution domains, outperforming both LLM-simulated training and concurrent synthesis methods. We believe AWM’s 1,000 synthesized environments and its scalable pipeline are valuable resources to the community.
@@ Acknowledgment
We would like to thank Jeff Rasley for valuable discussions and for helping organize relevant resources.
@@ A.5 Training and Evaluation Details
Training Hyperparameters. Table 8 summarizes the hyperparameters used for GRPO training (Shao et al., 2024). According to the common agentic RL training settings (Li et al., 2025b; Wang et al., 2025b), we set the KL coefficient to 0.001. Beyond standard GRPO settings, we use a higher clip ratio of 0.28, following the recommendation in DAPO (Yu et al., 2025) to allow more exploration for the agent. Also, we use sequence-level importance sampling to mitigate the distribution shift issue between the rollout engine and the model training engine (Yao et al., 2025).
@@ B.1 Analysis on Step-level Format Correctness Reward
B.2 Case Study for Verification
@@
We provide three representative cases in Figures 28, 29, and 30 to illustrate why our code-augmented LLM-as-a-Judge is more robust than either rigid code-only verification or LLM-only judging from agent trajectories. For the first case (Fig. 28), the task is a clean, database-grounded query: the agent calls the correct tool to retrieve the full bid history of the item and summarize the findings; the verifier can deterministically confirm these statistics from the underlying state/returned records, and the judge aligns, illustrating the ideal regime where structured verification evidence is decisive. For the second case (Fig. 29), the environment exhibits an imperfection when the agent tries to create a routine; a strict code-only verifier that expects a state delta would incorrectly flag failure because the initial and final snapshots appear identical, yet the trajectory shows an idempotent success path, and the code-augmented judge uses this context to correctly mark Completed, reducing false negatives under transient tool/infrastructure issues. For the last case (Fig. 30), an API/tool calling error misleads the agent into creating a duplicate event and then adding the session under the wrong event ID; tool calls succeed locally so a judge without verifier grounding could be fooled into marking Completed, but the verifier reveals the real target event remains unchanged, enabling the code-augmented judge to identify a wrong operation and correctly reject the spurious success claim.

### v3

@@ 6.2 Analysis on Verification Design
Table 10: Code-augmented LLM-as-a-Judge reliability on 100 sampled trajectories evaluated 5 times each. Here we report the binary classification results (Completed vs. others).
Metric GPT-5.1 Sonnet-4.5 Qwen3.5
Self-Consistency ($\uparrow$) 90.8% 82.0% 76.3%
Pairwise Agreement ($\uparrow$) 95.5% 91.8% 88.1%
Fleiss’ $\kappa$ ($\uparrow$) 0.891 0.826 0.728
Reward Flip Rate ($\downarrow$) 9.2% 18.0% 23.7%
Table 9 compares three verification strategies. LLM-only verification is ungrounded in state changes and yields the weakest performance. Code-only verification improves over it but is brittle under environment imperfections, which can produce false negatives. Our code-augmented strategy combines structured state-diff signals with a reasoning LLM, achieving the best results across all model scales and benchmarks. The extra cost of GPT-5 as the judge is about \$1.80 per training step (at most 1,024 samples), and the asynchronous setting introduces negligible latency. We also study the format correctness reward in Appendix B.1.
Judge reliability. To verify the judge reliability for RL, we sample 100 trajectories and judge each 5 times with three different LLMs in Table 10. GPT-5.1 attains 95.5% pairwise agreement and Fleiss’ $\kappa$ of 0.891 with only a 9.2% reward flip rate, justifying the high reliability for RL training. Even open-source Qwen3.5-122B-A10B model reaches 88.1% agreement, and the high cross-model agreement indicates that our designed verification rubrics and signals capture genuine task completion rather than model-specific ones. More details are provided in Appendix B.2.
Figure 4: Scaling curve of AWM over environment quantity with the 4B model. Hatched bars denote EnvScaler which synthesized 191 environments for SFT and RL training.
@@ 6.3 Analysis on History-Aware Training
Table 11 compares our history-aware training against a full-history variant (w/o HL) under aligned and misaligned inference. When aligned, AWM w/ HL achieves the best results among all settings, indicating the benefit of optimizing directly under truncated histories. In contrast, the full-history variant is relatively insensitive to misalignment. When truncated at inference, it even shows a slight improvement in $\tau^{2}$, consistent with truncation suppressing interference from irrelevant turns. These findings suggest that agent history context management should be part of policy optimization rather than a purely inference-time heuristic.
@@ 6.4 Environment Scaling Curve
Figure 4 studies the scaling curve with respect to the number of training environments. Training on 10 environments leads to severe performance degradation, likely due to overfitting to the limited distribution. Scaling to 100 environments yields substantial gains, and at a matched count, AWM outperforms EnvScaler on both BFCLv3 and MCP-Universe benchmarks. Further scaling to 526 environments continues to improve agent performance. This monotonic improvement highlights the scalability of AWM for agentic RL.
@@ 7 Conclusion
In this paper, we propose Agent World Model (AWM), a scalable pipeline that synthesizes executable agentic environments by mirroring software development. Using AWM, we successfully scale to 1,000 environments with 10,000 tasks. These environments are code-driven and backed by SQL databases exposed through a unified MCP interface, supporting parallel isolated instances for large-scale agentic RL. Experiments on three benchmarks show that agents trained on our synthetic environments generalize well to out-of-distribution domains, outperforming both LLM-simulated training and concurrent synthesis methods. We believe AWM’s scalable pipeline and its 1,000 synthesized environments are valuable resources to the community.
@@
Robustness & Safety. AWM only exposes agents to noise arising from imperfect synthesis (74% of environments contain at least one bug, $\sim$4% of RL rollouts hit runtime errors), and does not evaluate behavior under intentionally injected adversarial perturbations such as corrupted database rows, malicious tool outputs, or prompt-injection content in observations. Since task generation assumes a post-authentication context for tractability, access-control behavior is also not directly trained, thus an over-eager agent could in principle perform unauthorized data access, irreversible state mutations, or unintended call chains. Some synthesized scenarios further touch sensitive domains (e.g., finance and healthcare) where errors carry asymmetric costs, despite AWM containing no real personal data. We therefore view adversarial robustness evaluation, access control training, and domain specific human-in-the-loop oversight as prerequisites for production deployment.
@@ Acknowledgment
We would like to thank Jeff Rasley for helping organize AWM’s related open-source resources.
@@ A.5 Training and Evaluation Details
Training Hyperparameters. Table 12 summarizes the hyperparameters used for GRPO training (Shao et al., 2024). According to the common agentic RL training settings (Li et al., 2025c; Wang et al., 2025b), we set the KL coefficient to 0.001. Beyond standard GRPO settings, we use a higher clip ratio of 0.28, following the recommendation in DAPO (Yu et al., 2025) to allow more exploration for the agent. Also, we use sequence-level importance sampling to mitigate the distribution shift issue between the rollout engine and the model training engine (Yao et al., 2025).
@@ B.1 Analysis on Step-level Format Correctness Reward
B.2 Detailed LLM-Judge Reliability
We elaborate on the reliability study summarized in Sec. 6.2. The full results are shown in Table 13. We sample 100 agent trajectories from the 8B agent during training and judge each trajectory 5 times with each of three judges (GPT-5.1, Claude-4.5-Sonnet, Qwen3.5-122B-A10B), using default or recommended temperature of each model. We compute four metrics: Self-Consistency (fraction of trajectories on which a single judge returns the same label in all 5 trials), Pairwise Agreement (fraction of pairs of trials that agree), Fleiss’ $\kappa$, and Reward Flip Rate (fraction of trajectories where at least one trial disagrees with the majority). Even under the strictest 4-class setting, GPT-5.1 attains 91.2% pairwise agreement, and the open-source Qwen3.5 reaches 82.7%. This high reliability validates the use of our code-augmented LLM-as-a-Judge for reward assignment during RL training.
Table 13: Detailed LLM-judge reliability on 100 sampled trajectories evaluated 5 times each by three different judges. We report two settings: binary classification (Completed vs. others) which is the actual RL reward signal, and detailed 4-class classification across all reward labels (Completed / Partially Completed / Agent Error / Environment Error).
Metric GPT-5.1 Claude-4.5-Sonnet Qwen3.5-122B-A10B
Binary (Completed vs. others)
Self-Consistency ($\uparrow$) 90.8% 82.0% 76.3%
Pairwise Agreement ($\uparrow$) 95.5% 91.8% 88.1%
Fleiss’ $\kappa$ ($\uparrow$) 0.891 0.826 0.728
Reward Flip Rate ($\downarrow$) 9.2% 18.0% 23.7%
4-class (full reward classification)
Self-Consistency ($\uparrow$) 82.7% 65.0% 69.5%
Pairwise Agreement ($\uparrow$) 91.2% 83.5% 82.7%
Fleiss’ $\kappa$ ($\uparrow$) 0.781 0.693 0.650
Reward Flip Rate ($\downarrow$) 11.2% 25.0% 18.6%
B.3 Code-Level Diversity Analysis
To complement the embedding-based diversity analysis in Figure 3, we measure cross-environment similarity directly at the code level for all 1,000 environments in Table 14. We use the AST-based Tree-Structure Edit Distance (TSED) implementation [^2] https://github.com/mizchi/similarity to detect duplicated function bodies across files, and token-level / endpoint-name / class-name Jaccard for lexical comparison. The results show that no function-level structural duplicates exist across any pair of environments at TSED threshold 0.5, endpoint and class names have cross-environment Jaccard well below 0.01, and token Jaccard stays low (max 0.324 and the residual may come from shared Python/FastAPI keywords). The code-level evidence corroborates our embedding-level diversity finding: AWM pipeline is suitable for large-scale agent environments synthesis.
Table 14: Code-level similarity across AWM environments. We measure structural similarity via AST-based Tree-Structure Edit Distance (TSED), and lexical similarity via token / endpoint / class-name Jaccard.
Metric Granularity Mean Similarity Max
AST Function Duplicate (TSED, threshold $\geq 0.5$) function-level 0.0% 0.0%
Token Jaccard token-level 0.183 0.324
Endpoint Name Jaccard API route-level 0.004
Class Name Jaccard schema-level 0.009
B.4 Case Study for Verification
@@
We provide three representative cases in Figures 28, 29, and 30 to illustrate why our code-augmented LLM-as-a-Judge is more robust than either rigid code-only verification or LLM-only judging from agent trajectories. For the first case (Fig. 28), the task is a clean, database-grounded query: the agent calls the correct tool to retrieve the full bid history of the item and summarize the findings; the verifier can deterministically confirm these statistics from the underlying state/returned records, and the judge aligns, illustrating the ideal regime where structured verification evidence is decisive. For the second case (Fig. 29), the environment exhibits an imperfection when the agent tries to create a routine; a strict code-only verifier that expects a state delta would incorrectly flag failure because the initial and final snapshots appear identical, yet the trajectory shows an idempotent success path, and the code-augmented judge uses this context to correctly mark Completed, reducing false negatives under transient tool/infrastructure issues. For the last case (Fig. 30), an API/tool calling error misleads the agent into creating a duplicate event and then adding the session under the wrong event ID; tool calls succeed locally so a judge without verifier grounding could be fooled into marking Completed, but the verifier reveals the real target event remains unchanged, enabling the code-augmented judge to identify such wrong operation.